Consider a demand curve given by Q = 10.00 - (0.20)P. What is the inverse demand curve associated with this demand curve? ChoHow many units would the consumer demand if the good were free?
Answer #1

Answer
An inverse demand curve is
Q=10-0.2P
0.2P=10-Q
P=50-5Q
the inverse demand curve is
P=50-5Q

Option A
=======
Part 2
Maximum price consumer would pay for the first good is
P=50-5*1=$45
====
If the good is free then the consumer will buy
Q=10-0.2*0=10 units
the consumer will consume 10 units if it is free.